OpenStudy (anonymous):

Suppose the line tangent to the graph of f at x=3 is y=2x+1and suppose y=3x-3 is the line tangent to the graph of g at x=3. Find an equation of the tangent line to the following curves at x=3. a. f(x) + g(x) b. f(x) - 3g(x) c. 4f(x)

OpenStudy (anonymous):

NEED HELP!

OpenStudy (anonymous):

ok you know that \[f(3)=7,g(3)=6, f(3)+g(3)=13\]and you know that the slope will be 2 + 3 = 5 so you can use the point- slope formula to find the equation

OpenStudy (anonymous):

its it clear how you know these things?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

ok then are you good from there? because the others work the same way. you know \[(f+g)'=f'+g'\] \[(f+g)'(3)=f'(3)+g'(3)=3+2=5\] and \[(f+g)(3)=6+7=13\] so use \[y-13=5(x-3)\] for your first answer

OpenStudy (anonymous):

i think i might be good for the rest, thank you

OpenStudy (anonymous):

i got 5x+3 for b but thats not right. could you show me the steps and the answesr cause im not sure what i did wrong